OpenStudy (anonymous):

Find the derivative of f(x) = -10/x at x = -12.

OpenStudy (anonymous):

\[f \prime (x) = \frac{ 10 }{ x ^{2} }\] \[f \prime (-12) = \frac{ 10 }{ (-12)^{2}} = \frac{ 10 }{ 144 } = \frac{ 5 }{ 72 }\]